OpenStudy (anonymous):

The average marks on a test taken by a group of 9 students was 59. Jack was absent . What will he need to score to make the average group 60

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (butterflydreamer):

okay firstly, we know that the average for 9 students was 59. Work out their total marks by multiplying 9 and 59 = ?

OpenStudy (butterflydreamer):

sooo 9 *59 = ? :)

OpenStudy (anonymous):

531

OpenStudy (butterflydreamer):

yes! Okay now, we know that Jack was absent.. So if Jack is PRESENT this will mean there will be 10 people in total right? We want the average of these 10 students to equal to 60. So, work out their total scores.... 10 * 60 = ?

OpenStudy (anonymous):

600

OpenStudy (butterflydreamer):

Yeep, so Jack will need to score 600 - 531 = ?

OpenStudy (anonymous):

69

OpenStudy (butterflydreamer):

correct! :D

OpenStudy (anonymous):

thankss

OpenStudy (butterflydreamer):

you're welcomeeee